OpenStudy (anonymous):

Patterns, Function and Algebra Sue had 700 dollars in the bank on January 1. She deposits 30 dollars every week. Tom has 900 dollars in the bank and withdraws 20 dollars every week. In how many weeks will Sue’s amount of money in the bank exceed Tom’s? I am lost how to start this function, so that I can solve this. If anyone can just start it for me, I would greatly appreciate it.

OpenStudy (anonymous):

Okay this is what you have to do

OpenStudy (anonymous):

you know Sue adds 30 every week

OpenStudy (anonymous):

and she has 700

OpenStudy (anonymous):

Tom has 900 and he loses 20 a week

OpenStudy (anonymous):

So make a table

OpenStudy (anonymous):

k, so I can use function equation for it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=20x+900

OpenStudy (anonymous):

that's where im having the brain fart

OpenStudy (anonymous):

my bad -20x

OpenStudy (anonymous):

y=-20x+900

OpenStudy (anonymous):

y=30x+700

OpenStudy (anonymous):

and then y=30x+700

OpenStudy (anonymous):

just graph it and use slope-intercept form

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You know what slope-intercept form is don't you?

OpenStudy (anonymous):

ok, thanks a bunch.. my brain is fried and I think it was so simple I couldn't think of it. lol

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

haha you're welcome

OpenStudy (anonymous):

Tell me if you need any further assistance in Algebra 1 took it last year

OpenStudy (anonymous):

taking Geometry right now

OpenStudy (anonymous):

I'm taking geometry this year too..I'm switching from home-school to a catholic high school for my 9th grade year. I'm just finishing up my algebra I, so I am ready for geometry in 2 weeks

OpenStudy (anonymous):

Oh cool

OpenStudy (anonymous):

I am going to take Algebra 2 this upcoming school year

OpenStudy (anonymous):

Taking Geometry on FLVS

OpenStudy (anonymous):

(Florida Virtual School)

OpenStudy (anonymous):

yeah, cant play football in high school unless I go to an physical school

OpenStudy (anonymous):

haha true

OpenStudy (anonymous):

cool used to go to treca digital academy in ohio

OpenStudy (anonymous):

Okay got to go

OpenStudy (anonymous):

Can you give me a medal?

OpenStudy (anonymous):

For helping you?

OpenStudy (anonymous):

medal given

OpenStudy (mathstudent55):

x = number of weeks. You have the two equations correct: Sue's money goes up as y = 700 + 30x Tom's money goes down as y = 900 - 20x Now set the equations equal and solve for x. 700 + 30x = 900 - 20x The answer will give you the number of weeks it takes for Sue and Tom ot have the same amount of money. When you add 1 week to that, you 'll have your answer.